Sue and Tim are married taxpayers in the 32% marginal tac bracket. In 2018, they sold common stock shares, which they held for more than 3 months, for a capital gain of $3,800. They also sold some technology stock for a short-term capital loss of $9,000. In addition, they sold the home they had lived in for the past 10 years and experienced a $75,000 gain on the house. How much will their met capital gain (or losses) be for 2018? How much will they pay (or save) in taxes as a result of these transactions?
Answer #1

Profit/Loss in share transaction= (3800-9000)=-$5200

As Sue and Tim was living for more than 2.5 years and they couple so they will be tax exempted up to $500000. So, they will not pay any tax on gain of $75000 from selling of house.

Net Capital Gain= (75000-5200)=$69800

They will not pay any taxes.